# Minimax Hypothesis Testing for the Bradley-Terry-Luce Model

### v1

@@ III-A Minimax Formulation and Decision Rule
$\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}(\mathbf{1}_{n}\pi^{\mathrm{T}})$ $\|^{2}_{\pi^{-1},\mathrm{F}}=\|\Pi P-P^{\mathrm{T}}\Pi\|^{2}_{\pi^{-1},\mathrm{F}}+\|P+P^{\mathrm{T}}-{\mathcal{P}_{\mathcal{E}}}(\mathbf{1}_{n}\mathbf{1}_{n}^{\mathrm{T}})\|^{2}_{\pi,\mathrm{F}}.$
The proof of Proposition 3 is provided in Section IV-C. (We note that since $\pi_{i}>0$ for all $i\in[n]$, the norm $\|\cdot\|_{\pi^{-1},\mathrm{F}}$ is well-defined.) Using Proposition 3, we show in Theorem 1 that if a pairwise comparison model $P$ satisfies Assumptions 1, 2, 3 and 4, then the quantity $\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}\pi^{\mathrm{T}})\|_{\mathrm{F}}/\|\pi\|_{\infty}$ is always within constant factors of the Frobenius-norm-distance between $P$ and the set of BTL models (or more precisely, the set $\mathsf{BTL}_{h}$ defined below). Hence, $\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}\pi^{\mathrm{T}})\|_{\mathrm{F}}/\|\pi\|_{\infty}$ captures a natural notion of separation distance in this setting.
@@
Suppose the pairwise comparison matrix $P$ and the induced graph ${\mathcal{G}}$ satisfy Assumptions 1, 2, 3 and 4. Then, there exist constants $c_{1},c_{2}>0$ (independent of $n$) such that
$c_{1}$ $\frac{\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}\pi^{\mathrm{T}})\|_{\mathrm{F}}}{\|\pi\|_{\infty}}\leq\min_{B\in\mathsf{BTL}_{h}}\|P-{\mathcal{P}_{\mathcal{E}}}(B)\|_{\mathrm{F}}\leq c_{2}\frac{\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}\pi^{\mathrm{T}})\|_{\mathrm{F}}}{\|\pi\|_{\infty}},$
where $\mathsf{BTL}_{h}$ is the set of all pairwise comparison matrices $B$ corresponding to BTL models whose skill scores $\alpha\in{\mathbb{R}}^{n}$ satisfy Eq. 18 with $h=\delta^{-1}$.
The proof is provided in Section IV-D. It utilizes a new Lagrangian and perturbation-based approach to compute the Frobenius-norm-distance between a given pairwise comparison matrix and its closest reversible counterpart. This approach may be of independent utility in other matrix theoretic scenarios. We also note that our approach of measuring separation distance is quite different to distance measures between Markov chains based on spectral radius introduced in [67, 68]. Next, armed with Proposition 2 and Theorem 1, we formally define the hypothesis testing problem for BTL models.
@@ III-A1 Hypothesis Testing Problem
For a given tolerance parameter $\epsilon>0$ and a pairwise comparison model $P$ satisfying Assumptions 1, 2, 3 and 4, we can formulate the hypothesis testing problem in (5) as:
@@ III-A2 Minimax Risk and Decision Rule
Let $\Phi$ denote a decision rule or hypothesis test that maps the consolidated observations ${\mathcal{Z}}$ to $\{0,1\}$, where 00 represents the null hypothesis $H_{0}$ and $1$ represents the alternative hypothesis $H_{1}$. Let ${\mathbb{P}}_{H_{0}}$ and ${\mathbb{P}}_{H_{1}}$ denote the probability distributions of the observations ${\mathcal{Z}}$ under $H_{0}$ and $H_{1}$, respectively. Throughout this paper, we will use ${\mathbb{P}}_{H_{0}}$ and ${\mathbb{P}}_{H_{1}}$ when we need to specify a hypothesis, and ${\mathbb{P}}$ when a probability expression holds for both hypotheses. Similarly, ${\mathbb{E}}_{H_{0}}$, ${\mathbb{E}}_{H_{1}}$, and ${\mathbb{E}}$ will represent the corresponding expectation operators. Furthermore, for a fixed induced graph ${\mathcal{G}}$ satisfying Assumption 3, let ${\mathcal{M}}_{0}$ and ${\mathcal{M}}(\epsilon)$ denote the sets of pairwise comparison models that satisfy the null and alternative hypotheses, respectively, in (19) along with Assumptions 1, 2 and 4:
@@
Now, we can define the minimax risk for the graph ${\mathcal{G}}$ as
$\mathcal{R}_{\mathsf{m}}[{\mathcal{G}}]\triangleq\inf_{\Phi}\left\{\sup_{P\in{\mathcal{M}}_{0}}{\mathbb{P}}_{H_{0}}(\Phi({\mathcal{Z}})=1)+\sup_{P\in{\mathcal{M}}(\epsilon)}{\mathbb{P}}_{H_{1}}(\Phi({\mathcal{Z}})=0)\right\},$ (22)
@@
$\varepsilon_{\mathsf{c}}=\inf\left\{\epsilon>0:\mathcal{R}_{\mathsf{m}}[{\mathcal{G}}]\leq\frac{1}{2}\right\}.$ (23)
The choice of constant $\frac{1}{2}$ is arbitrary and could be replaced by any constant in $(0,{1})$.
Formally, one of our primary goals is to provide bounds on the critical threshold and determine its scaling with problem parameters like $n$. Intuitively, when $\epsilon$ is larger than $\varepsilon_{\mathsf{c}}$ (in scaling), the minimax risk can be made arbitrarily small, but if $\epsilon$ is smaller than $\varepsilon_{\mathsf{c}}$, then the minimax risk cannot be made small. To analyze the critical threshold, we introduce a statistical test that takes the consolidated observations ${\mathcal{Z}}$ as input and thresholds the following statistic:
$T\triangleq\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}\bigg{(}\left(\hat{\pi}_{i}+\hat{\pi}_{j}\right)^{2}\frac{Z_{ij}(Z_{ij}-1)}{k_{ij}(k_{ij}-1)}+\hat{\pi}_{j}^{2}-2\hat{\pi}_{j}\left(\hat{\pi}_{i}+\hat{\pi}_{j}\right)\frac{Z_{ij}}{k_{ij}}\bigg{)}{\mathbbm{1}}_{k_{ij}>1},$ (24)
@@
${\hat{S}}_{ij}\triangleq\begin{cases}\frac{Z_{ij}}{k_{ij}d},&(i,j)\in{\mathcal{E}}\text{ and }i\neq j\\ 1-\frac{1}{d}\sum\limits_{\begin{subarray}{c}u\in[n]\backslash\!\{i\}:\\ (i,u)\in{\mathcal{E}}\end{subarray}}\frac{Z_{iu}}{k_{iu}},&i=j\\ 0,&\text{otherwise }\end{cases}.$ (25)
@@
Lastly, we remark that testing for a single BTL model can be conducted using a $\chi^{2}$-type goodness-of-fit test [69], and testing for a class of BTL models could be attempted using a composite $\chi^{2}$-test [70]. However, performing sharp minimax analysis on the corresponding $\chi^{2}$-statistics poses challenges. Therefore, our focus lies on the proposed test statistic $T$ in Eq. 24. Notably, we demonstrate minimax optimality of the critical threshold for complete graphs, thereby underscoring the effectiveness of considering our proposed test.
@@ III-B Upper Bound on Critical Threshold
Consider the hypothesis testing problem in (19) such that Assumptions 1, 2, 3 and 4 hold. Then, there exist constants $c_{3},c_{4},c_{5}>0$ such that if $n\geq c_{3}$, the number of comparisons per pair satisfies $k\geq\max\!\big{\{}2,\frac{c_{4}h\log n}{d_{\max}\xi^{4}}\big{\}}$, and $d_{\max}\geq(\log n)^{4}$, the critical threshold defined in Eq. 23 is upper bounded by
$\varepsilon_{\mathsf{c}}^{2}\leq\frac{c_{5}}{nk}.$
The proof is provided in Section V. The proof has several essential steps. Among these, the most important step involves computing $\ell^{2}$-error bounds (see Lemma 6) for the estimated BTL parameters $\hat{\pi}$ when the data is generated by a general pairwise comparison model, which is not necessarily BTL (i.e., under $H_{1}$). The derivation of these error bounds requires us to analyze the second largest singular values of DTMs corresponding to non-reversible Markov chains, particularly in the context of induced graphs that are not complete. Once we have established these $\ell^{2}$-error bounds, we can compute bounds on ${\mathbb{E}}[T]$ and ${\mathrm{var}}(T)$. This step involves mitigating the correlation between the terms $\hat{\pi}$ and $\{Z_{ij}\}_{(i,j)\in{\mathcal{E}}}$ in Eq. 24 as both of them share the same source of randomness. Broadly speaking, this is done by splitting the product of dependent terms into three parts (by utilizing the identity $\hat{a}\hat{b}=(\hat{a}-a)(\hat{b}-b)+(\hat{a}-a)b+a\hat{b}$, where $\hat{a}$ and $\hat{b}$ are correlated estimators of $a$ and $b$) and bounding each part using appropriate concentration inequalities. Moreover, in the case where the induced graph is complete, the analysis becomes considerably simpler. An instance of such simplification in the proofs arises through the utilization of bounds between contraction coefficients of $S$ (see Section V-A and Lemma 3).
@@
$\Phi({\mathcal{Z}})={\mathbbm{1}}_{T\geq t},$ (27)
where the precise threshold $t=\Theta((nk)^{-1})$ is given in Eq. 60. (In other words, this decision rule returns the alternative hypothesis if and only if $T\geq t$ for an appropriate threshold $t=\Theta((nk)^{-1})$.) In Section VII-A, we also present a permutation-test-based approach to obtain a non-explicit threshold for our test based on data. This approach can be more readily employed in simulations and works even when the $k_{ij}$’s are not all equal.
@@
Consider the hypothesis testing problem in (19) such that Assumptions 1, 2, 3 and 4 hold. Then, there exist constants $c_{3},c_{4},c_{5}>0$ such that if $n\geq c_{3}$, the number of comparisons per pair satisfies $k\geq\max\!\big{\{}2,\frac{c_{4}h\log n}{d_{\max}\xi^{4}}\big{\}}$, and $d_{\max}\geq\log n$, the critical threshold defined in Eq. 23 is upper bounded by
$\varepsilon_{\mathsf{c}}^{2}\leq\frac{c_{5}(\log n)^{1/2}}{nk}.$
The proof of Proposition 4 can be gleaned from the proof of Theorem 2 in Section V. In essence, the behavior in Proposition 4 stems from the fact that the proof of Lemma 7 in Section V-C1 relies on a concentration inequality (see Proposition 10) which, under the stronger assumption $d_{\max}\geq(\log n)^{4}$, allows us to avoid the poly-logarithmic factor in Proposition 4 when establishing Theorem 2. However, if a standard matrix Bernstein inequality were employed for concentration, the “special” constants $c_{\alpha}$ and $c_{\gamma}$, defined in Lemma 7, would scale as $(\log n)^{1/2}$. Then, the proof of Proposition 4 would follow by using essentially the same logic as the proof of Theorem 2 and observing that Eq. 63 yields an additional factor of $(\log n)^{1/2}$ in the scaling of $\varepsilon^{2}_{\mathsf{c}}$. Moreover, in this regime, our decision rule returns the alternative hypothesis if and only if $T\geq\frac{\gamma\sqrt{\log n}}{nk}$ for some appropriately chosen constant $\gamma$.
[…]
Consider the hypothesis testing problem in (19) such that Assumptions 1, 2, 3 and 4 hold, and suppose that $d_{\max}\geq(\log n)^{4}$ and there exists a constant $c_{4}>0$ such that the number of comparisons per pair satisfies $k\geq\max\!\big{\{}2,\frac{c_{4}h\log n}{d_{\max}\xi^{4}}\big{\}}$. Then, we have the following bounds:

### v2

@@ III-A Minimax Formulation and Decision Rule
$\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}(\mathbf{1}_{n}\pi^{\mathrm{T}})$ $\|^{2}_{\pi^{-1},\mathrm{F}}=\|\Pi P-P^{\mathrm{T}}\Pi\|^{2}_{\pi^{-1},\mathrm{F}}$
$+\|P+P^{\mathrm{T}}-{\mathcal{P}_{\mathcal{E}}}(\mathbf{1}_{n}\mathbf{1}_{n}^{\mathrm{T}})\|^{2}_{\pi,\mathrm{F}}.$
The proof of Proposition 3 is provided in Section IV-C. (We note that since $\pi_{i}>0$ for all $i\in[n]$, the norm $\|\cdot\|_{\pi^{-1},\mathrm{F}}$ is well-defined.) Using Proposition 3, we show in Theorem 1 that if a pairwise comparison model $P$ satisfies Assumptions 1 and 2 with $h,\xi=\Theta(1)$, then the quantity $\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}\pi^{\mathrm{T}})\|_{\mathrm{F}}/\|\pi\|_{\infty}$ is always within constant factors of the Frobenius-norm-distance between $P$ and the set of BTL models (or more precisely, the set $\mathsf{BTL}_{h}$ defined below). Hence, $\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}\pi^{\mathrm{T}})\|_{\mathrm{F}}/\|\pi\|_{\infty}$ captures a natural notion of separation distance in this setting.
@@
Suppose the pairwise comparison matrix $P$ and the induced graph $\mathcal{G}$} satisfy Assumptions 1 and 2. Then, there exist constants $c_{1},c_{2}>0$ (independent of $n$) such that
$c_{1}\frac{\xi^{2}}{h^{3}}$ $\frac{\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}\pi^{\mathrm{T}})\|_{\mathrm{F}}}{\|\pi\|_{\infty}}\leq\min_{B\in\mathsf{BTL}_{h}}\|P-{\mathcal{P}_{\mathcal{E}}}(B)\|_{\mathrm{F}}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \leq c_{2}h\frac{\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}\pi^{\mathrm{T}})\|_{\mathrm{F}}}{\|\pi\|_{\infty}},$
where $\mathsf{BTL}_{h}$ is the set of all pairwise comparison matrices $B$ corresponding to BTL models whose skill scores $\alpha\in{\mathbb{R}}^{n}$ satisfy $\frac{\max_{i\in[n]}\alpha_{i}}{\min_{j\in[n]}\alpha_{j}}\leq h$.
The proof is provided in Section IV-D. It utilizes a new Lagrangian and perturbation-based approach to compute the Frobenius-norm-distance between a given pairwise comparison matrix and its closest reversible counterpart. This approach may be of independent utility in other matrix-theoretic scenarios. We also note that our approach of measuring separation distance is quite different to distance measures between Markov chains based on spectral radius introduced in [daskalakisDikkalaGravin2018, friedWolfer2022]. Next, armed with Proposition 2 and Theorem 1, we formally define the hypothesis testing problem for BTL models.
@@ III-A1 Hypothesis Testing Problem
For a given tolerance parameter $\epsilon>0$ and a pairwise comparison model $P$ satisfying Assumptions 1 and 2, we can formulate the hypothesis testing problem in (5) as:
@@ III-A2 Minimax Risk and Decision Rule
Let $\Phi$ denote a decision rule or hypothesis test that maps the consolidated observations $\mathcal{Z}$} to $\{0,1\}$, where 0 represents the null hypothesis $H_{0}$ and $1$ represents the alternative hypothesis $H_{1}$. Let ${\mathbb{P}}_{H_{0}}$ and ${\mathbb{P}}_{H_{1}}$ denote the probability distributions of the observations $\mathcal{Z}$} under $H_{0}$ and $H_{1}$, respectively. Throughout this paper, we will use ${\mathbb{P}}_{H_{0}}$ and ${\mathbb{P}}_{H_{1}}$ when we need to specify a hypothesis, and $\mathbb{P}$} when a probability expression holds for both hypotheses. Similarly, ${\mathbb{E}}_{H_{0}}$, ${\mathbb{E}}_{H_{1}}$, and $\mathbb{E}$} will represent the corresponding expectation operators. Furthermore, for a fixed induced graph $\mathcal{G}$}, let ${\mathcal{M}}_{0}$ and ${\mathcal{M}}(\epsilon)$ denote the sets of pairwise comparison models that satisfy the null and alternative hypotheses, respectively, in (17) along with Assumptions 1 and 2:
@@
Now, we can define the minimax risk for the graph $\mathcal{G}$} and a given $\epsilon>0$ as
$\mathcal{R}_{\mathsf{m}}[{\mathcal{G}},\epsilon]\triangleq\inf_{\Phi}\bigg\{$ $\sup_{P\in{\mathcal{M}}_{0}}{\mathbb{P}}_{H_{0}}(\Phi({\mathcal{Z}})=1)$
$\ \ \ \ \ +\sup_{P\in{\mathcal{M}}(\epsilon)}{\mathbb{P}}_{H_{1}}(\Phi({\mathcal{Z}})=0)\bigg\},$ (20)
@@
$\varepsilon_{\mathsf{c}}=\inf\left\{\epsilon>0:\mathcal{R}_{\mathsf{m}}[{\mathcal{G}},\epsilon]\leq\frac{1}{2}\right\}.$ (21)
The choice of constant $\frac{1}{2}$ is arbitrary and could be replaced by any constant in (0,1)(0,{1}).
Formally, one of our primary goals is to provide bounds on the critical threshold and determine its scaling with problem parameters like $n$. Intuitively, when $\epsilon$ is larger than $\varepsilon_{\mathsf{c}}$ (in scaling), the minimax risk can be made small, but if $\epsilon$ is smaller than $\varepsilon_{\mathsf{c}}$, then the minimax risk cannot be made small. To analyze the critical threshold, we introduce a statistical test that takes the consolidated observations $\mathcal{Z}$} as input and thresholds the following statistic:
$T\triangleq\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}\bigg(\left(\hat{\pi}_{i}+\hat{\pi}_{j}\right)^{2}$ $\frac{Z_{ij}(Z_{ij}-1)}{k_{ij}(k_{ij}-1)}+\hat{\pi}_{j}^{2}$
$\ \ \ \ \ -2\hat{\pi}_{j}\left(\hat{\pi}_{i}+\hat{\pi}_{j}\right)\frac{Z_{ij}}{k_{ij}}\bigg){\mathbbm{1}}_{k_{ij}>1},$ (22)
@@
[…]
@@
Lastly, we remark that testing for a single BTL model can be conducted using a $\chi^{2}$-type goodness-of-fit test [WassermanAllofStatistics], and testing for a class of BTL models could be attempted using a composite $\chi^{2}$-test [KAKIZAWAcompositechisquared]. However, performing sharp minimax analysis on the corresponding $\chi^{2}$-statistics poses challenges. Therefore, our focus lies on the proposed test statistic $T$ in Eq. 22. Notably, we demonstrate minimax optimality of the critical threshold for complete graphs, thereby underscoring the effectiveness of considering our proposed test.
@@ III-B Upper Bound on Critical Threshold
Consider the hypothesis testing problem in (17) such that Assumptions 1 and 2 hold. Then, there exist constants $c_{3},c_{4},c_{5}>0$ such that if $n\geq c_{3}$, the number of comparisons per pair satisfies $k\geq\max\!\big\{2,\frac{c_{4}h\log n}{d_{\max}\xi^{4}}\big\}$, and $d_{\max}\geq(\log n)^{4}$, the critical threshold defined in Eq. 21 is upper bounded by
$\varepsilon_{\mathsf{c}}^{2}\leq\left(\frac{c_{5}h^{3}}{\xi^{4}}\right)\frac{1}{nk}.$
The proof is provided in Section V. The proof has several essential steps. Among these, the most important step involves computing $\ell^{2}$-error bounds (see Lemma 6 and [Chenetal2019, Theorem 9]) for the estimated BTL parameters $\hat{\pi}$ when the data is generated by a general pairwise comparison model, which is not necessarily BTL (i.e., under $H_{1}$). The derivation of these error bounds requires us to analyze the second largest singular values of divergence transition matrices (DTMs) $R=\Pi^{1/2}S\Pi^{-1/2}$ [makur2019, MakurZheng2020, Huangetal2024], corresponding to non-reversible Markov chains $S$, particularly in the context of induced graphs that are not complete. Once we have established these $\ell^{2}$-error bounds, we can compute bounds on ${\mathbb{E}}[T]$ and ${\mathrm{var}}(T)$. This step involves mitigating the correlation between the terms $\hat{\pi}$ and $\{Z_{ij}\}_{(i,j)\in{\mathcal{E}}}$ in Eq. 22 as both of them share the same source of randomness. Broadly speaking, this is done by splitting the product of dependent terms into three parts (by utilizing the identity $\hat{a}\hat{b}=(\hat{a}-a)(\hat{b}-b)+(\hat{a}-a)b+a\hat{b}$, where $\hat{a}$ and $\hat{b}$ are correlated estimators of $a$ and $b$) and bounding each part using appropriate concentration inequalities. Moreover, in the case where the induced graph is complete, the analysis becomes considerably simpler. An instance of such simplification in the proofs arises through the utilization of bounds between contraction coefficients of $S$ (see Section V-A and Lemma 3).
@@
$\Phi({\mathcal{Z}})={\mathbbm{1}}_{T\geq{\mathsf{thr}}},$ (25)
where the precise threshold ${\mathsf{thr}}=\Theta(h^{3}/(nk\xi^{4}))$} is given in Eq. 57. (In other words, this decision rule returns the alternative hypothesis if and only if $T\geq{\mathsf{thr}}$ for an appropriate threshold ${\mathsf{thr}}=\Theta(h^{3}/(nk\xi^{4}))$.) In Section VII-A, we also present a permutation-test-based approach to obtain a non-explicit threshold for our test based on data. This approach can be more readily employed in simulations and works even when the $k_{ij}$’s are not all equal.
@@
Consider the hypothesis testing problem in (17) such that Assumptions 1 and 2 hold. Then, there exist constants ${\tilde{c}_{3},\tilde{c}_{4},\tilde{c}_{5}}>0$ such that if $n\geq{\tilde{c}_{3}}$, the number of comparisons per pair satisfies $k\geq\max\!\big\{2,\frac{{\tilde{c}_{4}}h\log n}{d_{\max}\xi^{4}}\big\}$, and $d_{\max}\geq\log n$, the critical threshold defined in Eq. 21 is upper bounded by
$\varepsilon_{\mathsf{c}}^{2}\leq\left(\frac{{\tilde{c}_{5}}h^{3}}{\xi^{4}}\right)\frac{(\log n)^{1/2}}{nk}.$
The proof of Proposition 4 can be gleaned from the proof of Theorem 2 in Section V. In essence, the behavior in Proposition 4 stems from the fact that the proof of Lemma 7 in Section V-C1 relies on a concentration inequality (see Proposition 9) which, under the stronger assumption $d_{\max}\geq(\log n)^{4}$, allows us to avoid the poly-logarithmic factor in Proposition 4 when establishing Theorem 2. However, if a standard matrix Bernstein inequality were employed for concentration, the “special” constants $c_{\alpha}$ and $c_{\gamma}$, defined in Lemma 7, would scale as $(\log n)^{1/2}$. Then, the proof of Proposition 4 would follow by using essentially the same logic as the proof of Theorem 2 and observing that Eq. 60 yields an additional factor of $(\log n)^{1/2}$ in the scaling of $\varepsilon^{2}_{\mathsf{c}}$. Moreover, in this regime, our decision rule returns the alternative hypothesis if and only if $T\geq\frac{\gamma h^{3}\sqrt{\log n}}{nk\xi^{4}}$ for some appropriately chosen constant $\gamma$.
We end this subsection with two remarks. Firstly, we can conclude by combining Theorems 1 and 2 that when the separation distance satisfies
$\min_{B\in\mathsf{BTL}_{h}}\|P-{\mathcal{P}_{\mathcal{E}}}(B)\|_{\mathrm{F}}\geq{\Omega}\bigg(\frac{h^{5/2}}{\sqrt{nk\xi^{4}}}\bigg),$ (26)
our test guarantees a minimax risk of at most $\frac{1}{2}$. This result demonstrates that the minimum separation distance required by our test increases with the principal ratio and when the underlying observation graph has “poor” expansion properties. Secondly, while it remains unclear whether the $k\geq 2$ condition is strictly necessary for the aforementioned results, we note that this condition was also required in a related two-sample testing problem [RastogiBalakrishnanShahAarti2022] where it was shown to be necessary. Intuitively, the $k\geq 2$ requirement arises because our test statistic is based on a second-moment (or variance) approximation, which requires at least two samples to obtain a meaningful estimate.
III-C Bounds on Type I and Type II Probabilities of Error
In this section, we provide bounds on the type I and type II probabilities of error for our proposed test. Here, the probability of type I error represents the probability of incorrectly rejecting the null hypothesis when it is true, while the probability of type II error represents the probability of failing to reject the null hypothesis when it is false.
Theorem 3 (Bounds on Type I and Type II Probabilities of Error).
Consider the hypothesis testing problem in (17) such that Assumptions 1 and 2 hold, and suppose that $d_{\max}\geq(\log n)^{4}$ and there exists a constant $c_{4}>0$ such that the number of comparisons per pair satisfies $k\geq\max\!\big\{2,\frac{c_{4}h\log n}{d_{\max}\xi^{4}}\big\}$. Then, we have the following bounds: